Random/ESG Robotics competition Rules

Preparation Rules: Spy reports may be provided by the class leaders. There is to be no other means of passing or stealing ideas between teams.

The Arena: The arena will be formed from a 4 ft x 8 ft piece of plywood and a 4 ft x 4 ft piece in the shape of an L. The arena will be bounded by 2x4s, on their 2-inch sides, placed on top of the plywood. Wood puddy will be used to smooth the crack between the sheets of plywood.

The Robots: Each team will have 3 robots.

At the start of the competitions, the robots may have at most a 1 ft x 1 ft square footprint. At no time during the competition may any appendage of the robot leave a 1.5 ft x 1.5 ft square, unless the appendage is connected only by nonstructural wires. No part of the robot may ever leave the arena (it may go on top of the wall, but not past it).

Each team may only use up to 4 handy boards between their 3 robots. No additional ports may be used unless the competitors build a interface with additional ports themselves. All power must come from the handy boards or nuclear reactors.

No robot evects may have a blast radius greater than 1 ft. No liquids may be used. IR may not be used to disrupt the flags’ signals. Damage to the arena will result in loss of the match. Damage to lego bricks and electrical components (except non-structural wires) will result in loss of the match. No robot effects may endanger people within 30 ft of the arena.

All permanent changes to legos must be cleared with the instructors. All nonstandard lego parts (i.e. non-legos) and devices must be cleared with the instructors. Small metal braces and perf-board may be used for structure. The only glue allowed is hot glue, and it may not be used to glue legos together.

The Flags: Each flag will be an IR beacon on a unique (known) frequency. These flags will be designed to be easily moved by either picking them up or pushing them. A team may not intentionally move its own flag. No robot may attempt to block the flags’ IR signals. At no point may a robot surround its own flag.

Competition Rules: The competition will take place over two days during the week before finals week. The first day’s competition will take place at ESG and the second day’s at Random. The teams will alternate on which side is to be near the crack between the plywood.

At the start of the competition, the flags will be in the center of the 2 ft x 4 ft homebase of each team. The robots may be positioned anywhere inside this region.

To win, one team must move the opponent’s flag from its opponent’s homebase to its own. The competition will continue until one side wins or the instructors stop the match. The instructors may stop the match when a penalty occurs or after 10 minutes. There may be no outside intervention during a match.

If a robot breaks a rule intentionally (the robot is so designed as to break the rule), that team loses the match. If it breaks a rule unintentionally twice, its team loses the match.